What is the lowest total resistance that can be secured by combinations of four coils of resistance 4 Ω , 8 Ω , 12 Ω , 24 Ω?

What is   the lowest total resistance that can be secured by combinations of four coils of resistance 4 Ω , 8 Ω , 12 Ω , 24 Ω?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


The lowest resistance is when the resistances are connected in parallel
Total resistance in parallel
= 1/R1 + 1/R2 + 1/R3 + 1/R4
1/R = 1/2 + 1/8 + 1/12 + 1/24
= 12/24
1/R = 1/2 ohm, R = 2 ohm
Thus, lowest resistance is 2 ohm.

Thanks